John is a 23 year old 58 kg male currently intubated and receiving mechanical ventilation. She is currently sedated on propofol at 20 mL/hr and receiving treatment for cryptococcal meningitis (secondary to drug abuse) with amphotericin B intravenously. Her past medical history is unremarkable. She was only able to tolerate about 15% of her meals (prior to intubation) due to the mental status changes related to the infection. You are asked to manage PPN and electrolytes on this patient.

  1. What would be the goal rate of a 4.25% amino acid/10% dextrose PPN formula if you wanted to provide 1.2 gm/kg of protein in John.?
  2. How many calories per day will John receive if the above rate of PPN solution is used and you added 250mL of 20% intralipids as a piggyback each day?
Answer #1

1. First of all calculate the amount of protein required by John

Weight of the patient is 58 kg

Amino acid % = 4.25

Protein = gm/ kg / day

= 1.2 gm× 58 kg

= 69.6gm / day

Gram per day ÷ ml / day = % of amino acid

= 69.6 ÷ ml / day = 4.25

69.6/4.25 = ml / day

= 16.37 ml / day is goal rate .

2. 1 gram of protein provides 4kcal

Then 69.6 gmof protein will provide = 69.6 × 4 = 278.4 kcal.

Plus addition of 250 ml of 20 % of intralipids will provide

2 kcal per ml = 250 ×2= 500ml

Hence ,the total amount of kcal received by John = 500 + 278.4 kcal

= 778.4 kcal each day